the income of a and b are in the ratio 3:2 ant there expenditure in the ratio 5:3. if each saves rs 1500 then b income ?

Answer by ankor@dixie-net.com(22740)

the income of a and b are in the ratio 3:2
let x = the multiplier
then
3x = a's income
and
2x = b's income
:
and their expenditures in the ratio 5:3.
let y = the multiplier
then
5y = a's expenditure
and
3y = b's expenditure
:
if each saves rs 1500 then b income ?
3x - 5y = 1500
and
2x - 3y = 1500
use elimination here. multiply the 1st eq by 3, the 2nd eq by 5
9x - 15y = 4500
10x - 15y =7500
-----------------Subtraction y find x
-1x = -3000
x = -3000/-1
x = +3000 is the multiplier
"then b income ?"
2(3000) = $6000 is b's income
:
:
You can confirm this yourself, find y and do the math